OpenStudy (sabrina1):

Suppose you have $2000 to invest. At what annual interest rate r (expressed in %) compounded continuously you have to invest in order to double your investment in 8 years ?

OpenStudy (anonymous):

2 = (1+r)^8 r = 2^(1/8) - 1 = 0.0905077 9.05%

OpenStudy (sabrina1):

thank you:)

OpenStudy (sabrina1):

i just have on question. why did you use 2

OpenStudy (anonymous):

because it's double

OpenStudy (sabrina1):

well i dont get to the same answer sorry. i think its because i have a diff formula thats why im confused. the formula im given is:A(t)=p*e^rt where e=2.72